Wrennie’s grandfather had left Wrennie a sum of RM83,000 for her undergraduate studies. The payment will be made available to her on her 18th birthday. Wrennie has just turned 6 years old. Before she turns 18 years old, her parents plan to deposit the sum of money with the OSK Trustee. Based on the past record, the OSK Trustee have been providing 4% per annum semi-annually. This rate of return is expected to continue for the foreseeable future. Wrennie’s parents plan to use the money inherited to cover the costs of education and her living expenses. Upon Wrennie turning 18 years old, the OSK Trustee provides a return of 8% per annum paid monthly for the proceeds which was invested in a managed fund earlier.

Required: i) Calculate the payment that Wrennie can expect to receive each month if she uses all the proceeds during the course of her five year of dentistry degree?
Answer #1

The annual rate of interest is 4% per annum, after dividing the annual rate with 2, the semi-annual rate of interest comes out to be 2%.

Compute the PVIF at 2% for 24 periods, using the equation as shown below:

PVIF = 1/ (1 + Rate)Number of periods

              = 1/ (1 + 0.02)24

         = 1/ 1.608437

         = 0.62172149

Hence, the PVIF at 2% and 24 periods is 0.62172149.

Compute the value of deposit after 24 periods, using the equation as shown below:

Deposit value = Deposit amount/ PVIF2%, 24

                       = RM83,000/ 0.62172149

                       = RM133,500.291731

Hence, the value of the deposit is RM133,500.291731

The annual rate is 8% after the age of 18 years. The monthly rate will be 0.6666667% which is calculated by dividing the 8% with 12 months.

Compute the PVIFA at 0.6666667% and 60 periods, using the equation as shown below:

PVIFA = {1 – (1 + Rate)-Number of periods}/ Rate

                   = {1 – (1 + 0.6666667%)-60}/ 0.6666667%

            = 49.3184

Hence, the PVIFA at 0.6666667% and 60 periods is 49.3184.

Compute the monthly payment, using the equation as shown below:

Monthly payment = Value of deposits/ PVIFA0.6666667%, 60

                                    = RM113,500.291731/ 49.3184

                              = RM2301.37822254

Hence, the monthly payment is RM2301.37822254